OpenStudy (anonymous):

make v the subject of the formula z = v -1 / v

OpenStudy (anonymous):

is it \[\large z=\frac{v-1}{v}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok how would you solve \[\frac{v-1}{v}=7\]?

OpenStudy (anonymous):

not sure. Z = v - 1 divided by v how do i get single v?

OpenStudy (anonymous):

lets solve \[\frac{v-1}{v}=7\] first step one, multiply both sides by \(v\) and get \[v-1=7v\]

OpenStudy (anonymous):

subtract \(v\) get \[-1=6v\] and so \[v=-\frac{1}{6}\]

OpenStudy (anonymous):

those steps clear? because it will be identical steps for \[\frac{v-1}{v}=z\] except for one step of combining like terms which is easy with \(7v-v\)

OpenStudy (anonymous):

so you make Z a number?

OpenStudy (anonymous):

that is what i did to explain the steps your answer will have a z in it

OpenStudy (anonymous):

thank you salellite! that makes sense now.

OpenStudy (anonymous):

yw hope it is more clear than before it is still a bit complicated when you have to factor out a \(v\) to get your answer

OpenStudy (anonymous):

could you help me solve 1 more question? john is twice as old as mary. 5 years ago he was 3 times as old. what is marys age now. i understand that J = 2M and J-5 = 3(m-5) but how to i factor out J/

OpenStudy (anonymous):

you can replace \(j\) by \(2m\) and solve \[2m-5=3(m-5)\]and now there is no \(j\) in it

OpenStudy (anonymous):

Oh! I see now. Cheers!

OpenStudy (anonymous):

cheers!